OpenStudy (anonymous):

A block of mass 3 kg slides down an inclined plane inclined at 30° with respect to horizontal. The frictional force between the block and the plane is 8.7 N and the length of the plane is 1 m. What is the work done by the net force when the block reaches the bottom of the plane? and why?

random231 (random231):

find the net force first

OpenStudy (anonymous):

is the answer 6j @random231

random231 (random231):

howd u get it?

OpenStudy (anonymous):

the net force is the force of gravity minus the force of friction Fnet = Fgravity - Friction Fnet = m*g*sin30 - 8.7 we use m*g*sin30 because only a component of the force of gravity is causing the block to slide down the incline. so Fnet = 6 Work = Fx where F is the net force and x is the distance W = 6 * 1 W = 6J

random231 (random231):

yah well actually 6.3 ok

random231 (random231):

u r right

OpenStudy (anonymous):

how did you do it?

random231 (random231):

same way theres no other alternative

OpenStudy (anonymous):

thanks!